OpenStudy (thefluffmuncher):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7).

OpenStudy (noseboy908):

Do you know the formula for slope given two points?

OpenStudy (noseboy908):

oh, uh, is that your answer?

OpenStudy (thefluffmuncher):

no this are the options

OpenStudy (noseboy908):

so, do you know the formula for slope given two points?

OpenStudy (thefluffmuncher):

no

OpenStudy (noseboy908):

ok, here it is: \[(y2-y1)/(x2-x1)\]

OpenStudy (noseboy908):

find the slope of your two points from it to start.

OpenStudy (the_fizicx99):

Ehh wrong. M = (y2-y1)/(x2-x1)

OpenStudy (the_fizicx99):

7-1/-4-5

OpenStudy (noseboy908):

that's what I put....

OpenStudy (the_fizicx99):

Nope you forgot the "m"

OpenStudy (noseboy908):

I figured it was assumed, considering I had put "the formula for slope from two given points" in the previous reply.

OpenStudy (the_fizicx99):

anyways just solve

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{-4}}\quad ,&{\color{blue}{ 7}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)